what is the bond type for ethylene oxide? use electronegativities to determine the type of bond. show your calculations. also specify if the bond is nonpolar covalent, polar covalent or ionic
1
Expert's answer
2019-02-28T04:03:19-0500

Ethylene oxide



Electronagativities:

χ(Η)= 2.20

χ(C)= 2.55

χ(O) = 3.44

Find ΔΧ(C-H) = 2.55 - 2.20 = 0.35 . ΔΧ(C-H) < 0.4 , the bond C-H is nonpolar covalent bond

Find ΔΧ(O-C) = 3.44-2.55 = 0.89. 0.4<ΔΧ(O-C)< 1.8 , the bond O-C is polar covalent

Find ΔΧ(C-C) = 2.55 - 2.55 =0 . ΔΧ(C-C) < 0.4 , the bond C-C in nonpolar covalent bond.